# HCMA-UNet: A Hybrid CNN-Mamba UNet with Inter-Slice Self-Attention for Efficient Breast Cancer Segmentation

### v1

@@ Abstract
Breast cancer lesion segmentation in DCE-MRI remains challenging due to heterogeneous tumor morphology and indistinct boundaries. To address these challenges, this study proposes a novel hybrid segmentation network, HCMA-UNet, for lesion segmentation of breast cancer. Our network consists of a lightweight CNN backbone and a Multi-view Inter-Slice Self-Attention Mamba (MISM) module. The MISM module integrates Visual State Space Block (VSSB) and Inter-Slice Self-Attention (ISSA) mechanism, effectively reducing parameters through Asymmetric Split Channel (ASC) strategy to achieve efficient tri-directional feature extraction. Our lightweight model achieves superior performance with 2.87M parameters and 126.44 GFLOPs. A Feature-guided Region-aware loss function (FRLoss) is proposed to enhance segmentation accuracy. Extensive experiments on one private and two public DCE-MRI breast cancer datasets demonstrate that our approach achieves state-of-the-art performance while maintaining computational efficiency. FRLoss also exhibits good cross-architecture generalization capabilities. The source code and dataset is available on this link.
@@
Medical Image Segmentation, DCE-MRI, Mamba, Inter-Slice Self-Attention, FRLoss
@@ I Introduction
Existing 3D breast cancer segmentation networks are primarily based on CNNs or Transformer architectures [6, 7, 8]. However, both face trade-offs between computational efficiency and the ability to capture local and global features necessary for accurate segmentation. To overcome CNNs’ local receptive field limitation of CNNs and the computational complexity of transformers, Mamba [9] is a novel architecture designed to achieve long sequence modeling with linear computational complexity by combining Gated Multi-Layer Perceptron (Gated MLP) with the H3 State Space Model.
Figure 1: Overview of HCMA-UNet. (a) Loss computation module, which integrates the conventional Dice and CE loss with our proposed Loss (FRLoss). (b) Output Block, which performs final classification based on the features from the last decoder layer. (c) Res Block for downsampling operations. (d) Dense Block for channel expansion.
@@
To address the limitation of unidirectional modeling in 2D image segmentation, VMamba [13] employs bidirectional state space models and positional embedding mechanisms. Building upon this design principle, we propose the Multi-view Inter-Slice Self-Attention Mamba (MISM) module to overcome the inherent limitations of existing 3D Mamba architectures. This module processes slices along three anatomical planes (sagittal, coronal, and axial) and models slice features using the Visual State Space Block (VSSB) from VMamba. Based on the intra-slice features extracted by VSSB, the module incorporates an Inter-Slice Self-Attention mechanism to enhance inter-slice feature correlations, achieving effective integration of both intra-slice and inter-slice information. By embedding this module into a lightweight convolutional network, we combine the local feature extraction capabilities of convolutions with the global modeling advantages of Mamba, enhancing feature learning capability while maintaining computational efficiency. The main contributions of our work are summarized as follows:
  • We propose a Hybrid CNN-Mamba with Inter-Slice Self-Attention UNet (HCMA-UNet), which integrates CNN’s local feature perception with Mamba’s superior sequential modeling capability for efficient 3D breast cancer segmentation in DCE-MRI.
@@
Figure 2: Overview of MISM and its sub-modules: ISSA and VSSB.
Figure 3: Cross-Scan Mechanism.
@@ II-B MISM
To enhance Mamba’s performance in 3D medical image segmentation, we propose the MISM. Our approach reslices 3D volumes into three orthogonal sets of 2D slices (axial, coronal, and sagittal), enabling comprehensive spatial correlation capture through multi-view analysis. The structure of MISM as illustrated in Fig. 2. MISM incorporates three synergistic components:(1) VSSB, which performs efficient bidirectional scanning operations along both the height and width dimensions of each 2D slice, effectively capturing the intra-slice features. (2) The ISSA mechanism models inter-slice dependencies by learning relationships between slices to preserve critical anatomical correlations across different layers. (3) The ASC strategy is proposed to address the computational redundancy of processing identical feature maps in three directions. Given that the axial plane simultaneously displays the complete structure and symmetrical features, we allocate 50% of the channels to axial plane learning, while equally distributing the remaining channels (25% each) to coronal and sagittal planes.
@@ II-B1 VSSB
II-B2 ISSA
To effectively capture the spatial dependencies between slices in medical volumes, we propose ISSA, a one-dimensional Self-Attention mechanism. Operating along the direction orthogonal to the VSSB scanning plane, ISSA achieves linear computational complexity. The query (Q), key (K), and value (V) features are generated through pointwise convolution from the input tensor $x\in\mathbb{R}^{B\times C\times L_{1}\times L_{2}\times L_{3}}$, where $B$ and $C$ denote batch size and number of channels respectively, and $L_{1}$, $L_{2}$, $L_{3}$ are the spatial dimensions. To capture inter-slice relation, we reshape these features along the $L_{3}$ direction (similar operations can be applied to the $L_{1}$ or $L_{2}$ direction) into the following dimensions:
@@ III-D Result Comparison
VSSB ASC ISSA Mamba3d Dice Prec Rec IoU VS MParams
@@ III-E Ablation experiment of MISM
To evaluate the performance of the HCMA-UNet network and the effectiveness of MISM comprehensively, we design six ablation experiments. The experiments primarily focus on three core components of MISM: VSSB, ASC strategy, and ISSA. The experimental results are listed in Table II. First, we construct a baseline network containing only convolutional layers. The experimental results demonstrate that as various components are gradually introduced, the network’s performance metrics exhibit a continuous upward trend. Among them, the complete HCMA-UNet architecture exhibits optimal performance. To compare with the direct application of bidirectional Mamba, we replace MISM with a 3DMamba block. Our method outperforms 3DMamba with improvements of 1.18%, 1.01%, and 2.63% in Dice, Precision, and Recall, respectively. These improvements demonstrate MISM’s effectiveness in medical images.

### v2

# HCMA-UNet: A Hybrid CNN-Mamba UNet with Axial Self-Attention for Efficient Breast Cancer Segmentation

@@ Abstract
[…]
@@
Medical Image Segmentation, DCE-MRI, Mamba, Axial Self-Attention, FRLoss
@@ I Introduction
Figure 1: Overview of HCMA-UNet. (a) Loss computation module that combines Dice and Ce losses with our proposed FRLoss. (b) Output Block performs final classification based on the features from the last decoder layer. (c) Res Block for downsampling operations. (d) Dense Block for channel expansion.
@@
Existing 3D breast cancer segmentation networks are primarily based on CNNs or Transformer architectures [6, 8]. However, both face trade-offs between computational efficiency and the ability to capture local and global features necessary for accurate segmentation. To overcome CNNs’ local receptive field limitation of CNNs and the computational complexity of transformers, Mamba [9] is a novel architecture designed to achieve long sequence modeling with linear computational complexity by combining Gated Multi-Layer Perceptron (Gated MLP) with the H3 State Space Model.
@@
To address the limitation of unidirectional modeling in 2D image segmentation, VMamba [13] employs bidirectional state space models and positional embedding mechanisms. Building upon this design principle, we propose the Multi-view Axial Self-Attention Mamba (MISM) module to overcome the inherent limitations of existing 3D Mamba architectures. This module processes slices along three anatomical planes (sagittal, coronal, and axial) and models slice features using the Visual State Space Block (VSSB) from VMamba. Based on the intra-slice features extracted by VSSB, the module incorporates an Axial Self-Attention mechanism to enhance inter-slice feature correlations, achieving effective integration of both intra-slice and inter-slice information. By embedding this module into a lightweight convolutional network, we combine the local feature extraction capabilities of convolutions with the global modeling advantages of Mamba, enhancing feature learning capability while maintaining computational efficiency. The main contributions of our work are summarized as follows:
Figure 2: Overview of MISM and its sub-modules: ASA and VSSB.
Figure 3: Cross-Scan Mechanism.
  • We propose a Hybrid CNN-Mamba with Axial Self-Attention UNet (HCMA-UNet), which integrates CNN’s local feature perception with Mamba’s superior sequential modeling capability for efficient 3D breast cancer segmentation in DCE-MRI.
@@ II-B MISM
To enhance Mamba’s performance in 3D medical image segmentation, we propose the MISM. Our approach reslices 3D volumes into three orthogonal sets of 2D slices (axial, coronal, and sagittal), enabling comprehensive spatial correlation capture through multi-view analysis. The structure of MISM as illustrated in Fig. 2. MISM incorporates three synergistic components:(1) VSSB, which performs efficient bidirectional scanning operations along both the height and width dimensions of each 2D slice, effectively capturing the intra-slice features. (2) The ASA mechanism models inter-slice dependencies by learning relationships between slices to preserve critical anatomical correlations across different layers. (3) The ASC strategy is proposed to address the computational redundancy of processing identical feature maps in three directions. Given that the axial plane simultaneously displays the complete structure and symmetrical features, we allocate 50% of the channels to axial plane learning, while equally distributing the remaining channels (25% each) to coronal and sagittal planes.
@@ II-B1 VSSB
II-B2 ASA
To effectively capture the spatial dependencies between slices in medical volumes, we propose ASA, a one-dimensional Self-Attention mechanism. Operating along the direction orthogonal to the VSSB scanning plane, ASA achieves linear computational complexity. The query (Q), key (K), and value (V) features are generated through pointwise convolution from the input tensor $x\in\mathbb{R}^{B\times C\times L_{1}\times L_{2}\times L_{3}}$, where $B$ and $C$ denote batch size and number of channels respectively, and $L_{1}$, $L_{2}$, $L_{3}$ are the spatial dimensions. To capture inter-slice relation, we reshape these features along the $L_{3}$ direction (similar operations can be applied to the $L_{1}$ or $L_{2}$ direction) into the following dimensions:
@@ III-D Result Comparison
VSSB ASC ASA Mamba3d Dice Prec Rec IoU VS MParams
@@ III-E Ablation experiment of MISM
To evaluate the performance of the HCMA-UNet network and the effectiveness of MISM comprehensively, we design six ablation experiments. The experiments primarily focus on three core components of MISM: VSSB, ASC strategy, and ASA. The experimental results are listed in Table II. First, we construct a baseline network containing only convolutional layers. The experimental results demonstrate that as various components are gradually introduced, the network’s performance metrics exhibit a continuous upward trend. Among them, the complete HCMA-UNet architecture exhibits optimal performance. To compare with the direct application of bidirectional Mamba, we replace MISM with a 3DMamba block. Our method outperforms 3DMamba with improvements of 1.18%, 1.01%, and 2.63% in Dice, Precision, and Recall, respectively. These improvements demonstrate MISM’s effectiveness in medical images.